# Exploration and Incentivized Participation in Randomized Trials

### v10

@@ Abstract
Participation incentives is a well-known issue inhibiting randomized controlled trials (RCTs) in medicine, as well as a potential cause of user dissatisfaction for RCTs in online platforms. We frame this issue as a non-standard exploration-exploitation tradeoff: an RCT would like to explore as uniformly as possible, whereas each “agent” (a patient or a user) prefers “exploitation”, i.e., treatments that seem best. We incentivize participation by leveraging information asymmetry between the trial and the agents. We measure statistical performance via worst-case estimation error under adversarially generated outcomes, a standard objective for RCTs. We obtain a near-optimal solution in terms of this objective: an incentive-compatible mechanism with a particular guarantee, and a nearly matching impossibility result for any incentive-compatible mechanism. We consider three model variants: homogeneous agents (of the same “type” comprising beliefs and preferences), heterogeneous agents, and an extension with estimated type frequencies.
Keywords: exploration-exploitation tradeoff, incentive-compatiblity, economic design, randomized controlled trials, clinical trials, multi-armed bandits
@@ 1 Introduction
Randomized Controlled Trials (RCTs) is a gold-standard way to evaluate safety and efficiency of medical treatments. A paradigmatic design considers $n=2$ alternatives, a.k.a. arms. [^2] While one arm is typically a “control group” (such as an old treatment or a placebo), this distinction is not essential throughout this paper. Each patient $t$ is assigned an arm $a_{t}$ independently and uniformly at random. An outcome $\omega_{t}$ is observed and evaluated according to some predefined metric $f(\cdot)$. The goal is to use the observations to estimate the counterfactual averages $\tfrac{1}{T}\,\sum_{t\in[T]}f(\omega_{a,t})$, $a\in[n]$, where $T$ is the number of patient. Independent per-patient randomization over arms is essential to remove selection bias and ensure statistical validity. Uniform randomization is usually preferred, as it minimizes the estimation error. In particular, one can obtain strong provable guarantees even if each arm’s outcomes are generated by an adversary, rather than sampled from some fixed (arm-dependent) distribution. [^3] In the same sense as “oblivious adversary” in adversarial bandits (Auer et al.,, 2002).
A well-known issue which inhibits RCTs is the difficulty of recruiting patients (e.g., see surveys Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). This issue may be especially damaging for large-scale RCTs for wide-spread medical conditions with relatively inexpensive treatments. Then finding suitable patients and providing them with appropriate treatments would be fairly realistic, but incentivizing patients to participate in sufficient numbers may be challenging.
Similar issues arise in user-facing online platforms, where experimentation is pervasive and RCTs, a.k.a. A/B testing, is a standard approach. [^4] For background on A/B testing, see e.g., Kohavi et al., (2009); Kohavi and Longbotham, (2017). While platform users are typically not informed of the particular experiments they are assigned to, the loss in performance (whether actual or perceived) may lead to user dissatisfaction. Having robust participation incentives would therefore be reassuring to the users, and may allow for more ambitious and risky experiments.
@@
We make two standard assumptions from economic theory: the RCT has the power to (i) prevent undesirable information disclosure to the agents, and (ii) commit to following a particular policy. These assumptions are realistic in clinical RCTs due to regulation. Indeed, the history of an ongoing RCT and the assigned treatment should not revealed to the patients (or their doctors) to the extent possible, to avoid biasing them in favor of better-performing treatments (Detry et al.,, 2012, pp. 26-30). [^6] In particular, different treatments are typically presented in the same way, e.g., as externally-similar pills. Moreover, the design of an RCT must be fixed in advance, spelling out all applicable decision points, and revealed to the patients (Arango et al.,, 2012). Experimentation in online platforms is typically not revealed to the users, and the commitment power is a standard assumption in the literature (e.g., see “incentivized exploration” in Related Work).
@@
While RCTs with $n=2$ arms is the paradigmatic case, our results seamlessly extend to $n>2$. Note that clinical RCTs with $\geq 3$ arms are studied in biostatistics, e.g., (Hellmich,, 2001; Freidlin et al.,, 2008), and are important in practice, e.g., (Parmar et al.,, 2014; Redig and Jänne,, 2015).
@@
Our modelling approach. We model “incentivized participation” as a mechanism design problem under a statistical objective and incentives constraints. An RCT (mechanism) sequentially interacts with self-interested agents like a multi-armed bandit algorithm: in each round, an agent arrives, the mechanism chooses an arm for this agent, and observes an outcome. The mechanism’s statistical objective is standard for RCTs: counterfactually estimate each arm’s performance on the whole population, even if the outcomes are generated by an adversary. [^7] In contrast, the standard objective for multi-armed bandits is to optimize cumulative payoffs. Formally, the mechanism optimizes the worst-case estimation error for these counterfactual estimates, denoted $\mathtt{ERR}$. If not for the incentives constraints, the mechanism would simply choose arms uniformly at random.
Agents have utilities over the outcomes, possibly different from the “societal value” estimated by the RCT. The RCT may also be interested in the empirical frequency of a particular outcome (we allow a generic formulation which captures both societal value and outcome frequency).
@@
Participation incentives are modeled as Bayesian Individual Rationality (BIR), a standard notion in economic theory. The mechanism commits to a some randomized policy that assigns an arm to each agent, depending on the history of interactions with the previous agents. The agent does not observe the history or the chosen arm (as discussed above). BIR states that each agent must prefer participation to her “outside option”, in expectation over her Bayesian beliefs.
While our statistical guarantees must be robust against adversarial outcomes, the outcome-generating process could be much more well-behaved. Crucially, the agents believe that it is. Specifically, the agents believe that one’s outcome distribution is determined by the assigned arm and one’s observable attributes (resp., medical history or user’s profile). While such beliefs may be incorrect, they appear necessary for tractability, and are particularly appropriate in clinical RCTs as a way to represent pre-existing medical knowledge.
@@
We consider three model variants. To isolate the core issues of incentivized participation, we consider on the fundamental case of homogeneous agents with the same beliefs and preferences. The general case of heterogeneous agents endows each agent with a type: observable attributes which determine the outcome distributions in the agent’s beliefs and unobservable (private) preferences over the outcomes which need to be elicited in an incentive-compatible way. [^8] In practice, preference elicitation would happen during registration/opt-in process. For online platforms, private preferences and their elicitation make sense only if the RCT has explicit opt-in. The latter is uncommon, but may be needed if the RCT involves high-stakes interactions (such as buying a house) or is potentially controversial. The third variant provides the mechanism with estimated type frequencies over the agent population; this allows to mitigate the harmful influence of rare-but-difficult agent types.
@@
Our results. Our mechanisms have simple two-stage designs: a relatively short warm-up stage of predetermined duration, followed by the main stage that needs to be counterfactually estimated. The data from the warm-up stage (warm-up data) creates information asymmetry in the main state which drives participation incentives, whereas the main stage determines the counterfactual estimates. A warm-up stage which influences the main stage is very common in practice. [^9] E.g., phase-I clinical trials to calibrate the dosage and “pilot experiments” in online platforms.
@@
Our guarantees focus on the main stage, given enough warm-up data. A relatively small amount of warm-up data, determined by the prior and the utility structure, suffices to bootstrap the main stage for an arbitrarily large number of agents. The warm-up data can be collected exogenously e.g., using paid volunteers, or endogenously by a mechanism in our model, e.g., via prior work on incentivized exploration (see “Related work” and Lemma 3.6).
@@
We obtain a near-optimal solution for each model variants described above, from homogeneous agents to heterogeneous agents to estimated type frequencies. First, we put forward a benchmark: a hypothetical policy that observes the true “state of the world” according to agents’ beliefs and optimizes a certain objective. Then, we design a mechanism whose estimation error ($\mathtt{ERR}$) is upper-bounded in terms of this benchmark. Finally, we prove that no mechanism achieves better estimation error, up to constant factors. Many ideas carry over from one model variant to the next.
@@
Our mechanisms have a desirable property that their “sampling distributions” (the distributions from which the arms are sampled) do not adapt to the data collected in the main stage. Indeed, clinical RCTs in practice typically do not adapt to data, except possibly at a few pre-determined “review points”,101010 This is because treating the patients and/or observing the outcomes may take a long time, and the doctors may find it difficult to coordinate across the patients. So, many/most patients may need to be treated in parallel. and neither do most RCTs in online platforms due to engineering considerations. In contrast, our negative results apply to a wide class of mechanisms, including the “few review points” mechanisms mentioned above.
@@
Our contribution is not the mechanisms per se, but the whole “journey”: from the new model to the benchmark to the matching upper/lower bounds, and from homogeneous agents to heterogeneous agents to estimated type frequencies. Conceptually, we connect the dynamic process and statistical estimators in the mechanism with a static information-design problem in the benchmark.
@@
We focus on patients’ reluctance to participate that stems from their preferences among the available treatments (as opposed to other factors, e.g., logistical difficulties). Further, this paper is only concerned with patients’ initial decision to participate in an RCT. Their incentives to remain in the trial, e.g., if the treatment appears ineffective, is outside of our scope. [^11] This is indeed an issue in practice, as per (FDA,, 2001, Section 2.1.7.2). However, it is unclear if any approach based on information asymmetry could possibly help here: once the patient learns about the actual outcome, the “principal” does not have much/any informational advantage left.
1.1 Related work
Incentivized exploration (IE) is a closely related model concerning experimentation on online platforms. Here the “platform” is a recommendation system which strives to incentivize users to explore when they would prefer to exploit. The platform issues recommendations (which the users can override), and makes them incentive-compatible via information asymmetry. Introduced in (Kremer et al.,, 2014; Che and Hörner,, 2018), IE is well-studied, see (Slivkins,, 2019, Chapter 11) for a survey. Most technically related are Mansour et al., (2020); Immorlica et al., (2019).
The main difference concerns performance objectives. IE strives to maximize the total reward under time-invariant outcome distribution, or just explore each arm once. Neither serves our statistical objective, which requires independent near-uniform randomization in each round: reward-maximization pulls a mechanism in a very different direction, and merely exploring each arm $N$ times does not ensure statistical validity under adversarially chosen outcomes. The economic constraints are different, too, because agents in our model can only choose whether to participate. [^12] This distinction is consequential only for $n>2$ arms. Then the economic constraint in IE is stronger than ours. Hence, our mechanisms, analyses and technical guarantees are very different from those in IE. However, our mechanisms reuse IE for the “warm-up stage” (see Lemma 3.6).
@@ 1.1 Related work
Economic design. Using information asymmetry to incentivize agents to take particular action is a common goal in information design (Bergemann and Morris,, 2019; Kamenica,, 2019). However, this literature tends to have a very different technical flavor. The key distinction is that designer’s utility is (typically) directly derived from each agent’s actions. In contrast, our statistical objective depends on the sampling distributions and does not add up across agents.
@@
Difficulties in patient recruitment are widely acknowledged as a significant barrier (e.g.,  Adams et al.,, 1997; Grill and Karlawish,, 2010). Its reasons are well-studied, e.g., see surveys (Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). Patients preferences among treatments are listed among the major factors, along with other factors such as insufficient trust, logistical difficulties, or aversion to randomization and loss of control. Prior work suggested increasing the probability of the “new” treatment as a possible remedy (Vozdolska et al.,, 2009; Jenkins and Fallowfield,, 2000; Pocock,, 1979; Karlawish et al.,, 2008), albeit without a formal model of participation incentives and without theoretical guarantees. These papers focus on the scenario when the new treatment appears ex-ante preferable compared to non-participation, and do not revise the treatment probabilities as the new data arrives. Other suggested remedies, focusing on eligibility criteria, patient outreach, trial logistics, etc., are not as relevant to this paper.
@@
Our mechanisms are superficially similar to patient preference trials (PPTs) (Torgerson and Sibbald,, 1998), in that both allow patient preferences to impact treatment allocation. The key difference is that PPTs allow the patients to directly choose the chosen treatment (potentially at the expense of the internal validity), whereas in our setting a patient’s report only affects the sampling distribution (hence the internal validity is not compromised). Another distinction is that we elicit patient preferences over the treatment outcomes, not over the treatments themselves.
@@
While our mechanism follows the protocol of multi-armed bandits (Slivkins,, 2019; Lattimore and Szepesvári,, 2020), it strives to choose arms uniformly at random, whereas a bandit algorithm strives to converge on the best arm. [^13] Under the standard objective of cumulative reward/loss optimization, as expressed by regret. More related is “pure exploration” (starting from Mannor and Tsitsiklis,, 2004; Even-Dar et al.,, 2006), where a bandit algorithm predicts the best arm under a stationary outcome distribution, after exploring “for free” (e.g., uniformly) for $T$ rounds. In contrast, our model allows adversarially chosen outcomes, and requires estimates for all arms and all outcomes. Note that bandit algorithms, whether regret-minimizing or “pure exploration”, are generally not geared to incentivize participation. Interestingly, both “stationary” and adversarially chosen outcome distributions – like, resp., agent beliefs and statistical objective in our model – are well-studied in bandits, as, resp., “stochastic” and “adversarial” bandits.
1.2 Map of the paper
We first present and analyze our model for the case of homogeneous agents (Sections 2 and 3). This is to introduce the core model and key ideas in a more accessible way, disentangling them from the additional subtlety and heavier notation needed for the general case. Heterogeneous agents are defined and treated in Section 4. Section 5 works out an extension with estimated type frequencies. The technical machinery in Sections 3, 4 and 5 is introduced gradually and heavily reused.

### v11

@@ Abstract
Participation incentives is a well-known issue inhibiting randomized controlled trials (RCTs) in medicine, as well as a potential cause of user dissatisfaction for RCTs in online platforms. We frame this issue as a non-standard exploration-exploitation tradeoff: an RCT would like to explore as uniformly as possible, whereas each “agent” (a patient or a user) prefers “exploitation”, i.e., treatments that seem best. We incentivize participation by leveraging information asymmetry between the trial and the agents. We measure statistical performance via worst-case estimation error under adversarially generated outcomes, a standard objective for RCTs. We obtain a near-optimal solution in terms of this objective: an incentive-compatible mechanism with a particular guarantee, and a nearly matching impossibility result for any incentive-compatible mechanism. We consider three model variants: homogeneous agents (of the same “type” comprising beliefs and preferences), heterogeneous agents, and an extension that leverages estimated type frequencies to mitigate the influence of rare-but-difficult agent types.
Keywords: exploration-exploitation tradeoff, incentive-compatibility, economic design, randomized controlled trials, clinical trials, multi-armed bandits
@@ 1 Introduction
Randomized Controlled Trials (RCTs) are gold-standard ways to evaluate the safety and efficacy of medical treatments. A paradigmatic design considers $n=2$ alternatives, a.k.a. arms. [^2] While one arm is typically a “control group” (such as an old treatment or a placebo), this distinction is not essential throughout this paper. Each patient $t$ is assigned an arm $a_{t}$ independently and uniformly at random. An outcome $\omega_{t}$ is observed and evaluated according to some predefined metric $f(\cdot)$. The goal is to use the observations to estimate the counterfactual averages $\tfrac{1}{T}\,\sum_{t\in[T]}f(\omega_{a,t})$, $a\in[n]$, where $T$ is the number of patients. Independent per-patient randomization over arms is essential to remove selection bias and ensure statistical validity. Uniform randomization is usually preferred, as it minimizes the estimation error. In particular, one can obtain strong provable guarantees even if each arm’s outcomes are generated by an adversary, rather than sampled from some fixed (arm-dependent) distribution. [^3] In the same sense as “oblivious adversary” in adversarial bandits (Auer et al.,, 2002).
A well-known issue that inhibits RCTs is the difficulty of recruiting patients (e.g., see surveys Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). This issue may be especially damaging for large-scale RCTs for widespread medical conditions with relatively inexpensive treatments. Then finding suitable patients and providing them with appropriate treatments would be fairly realistic, but incentivizing patients to participate in sufficient numbers may be challenging.
Similar issues arise in user-facing online platforms, where experimentation is pervasive and RCTs, a.k.a. A/B testing, are a standard approach. [^4] For background on A/B testing, see e.g., Kohavi et al., (2009); Kohavi and Longbotham, (2017). While platform users are typically not informed of the particular experiments they are assigned to, the loss in performance (whether actual or perceived) may lead to user dissatisfaction. Having robust participation incentives would therefore be reassuring to the users, and may allow for more ambitious and risky experiments.
@@
We make two standard assumptions from economic theory: the RCT has the power to (i) prevent undesirable information disclosure to the agents, and (ii) commit to following a particular policy. These assumptions are realistic in clinical RCTs due to regulation and standard practice. Indeed, the history of an ongoing RCT and the assigned treatment should not be revealed to the patients (or their doctors) to the extent possible, to avoid biasing them in favor of better-performing treatments (Detry et al.,, 2012, pp. 26-30). [^6] In particular, different treatments are typically presented in the same way, e.g., as externally-similar pills. Moreover, the design of an RCT must be fixed in advance, spelling out all applicable decision points, and revealed to the patients (Arango et al.,, 2012). For online platforms, the data from an ongoing RCT is typically not revealed to the users, and the commitment power is a standard assumption in the literature (see Related Work).
@@
While RCTs with $n=2$ arms are the paradigmatic case, our results seamlessly extend to $n>2$. Note that clinical RCTs with $\geq 3$ arms are studied in biostatistics, (e.g.,  Hellmich,, 2001; Freidlin et al.,, 2008), and are important in practice, (e.g.,  Parmar et al.,, 2014; Redig and Jänne,, 2015).
@@
Our model. We model “incentivized participation” as a mechanism design problem under a statistical objective and incentive constraints. An RCT (mechanism) sequentially interacts with self-interested agents like a multi-armed bandit algorithm: in each round, an agent arrives, the mechanism chooses an arm for this agent, and observes an outcome. The mechanism’s statistical objective is standard for RCTs: counterfactually estimate each arm’s performance on the whole population, even if the outcomes are generated by an adversary. [^7] In contrast, the standard objective for multi-armed bandits is to optimize cumulative payoffs. Formally, the mechanism optimizes the worst-case estimation error for these counterfactual estimates, denoted $\mathtt{ERR}$. If not for the incentive constraints for participation, the mechanism would simply choose arms independently and uniformly at random for each agent.
The role of $f$, the RCT’s metric, is as follows. Typically, $f(\omega)$ corresponds to the “societal value” of the outcome $\omega$. The RCT may also be interested in the empirical frequency of $\omega$, which is captured by setting $f={\bf 1}_{\left\{\,\omega\,\right\}}$, the indicator function that chooses this outcome. Our definition of $\mathtt{ERR}$ handles all metrics $f$ at once, and therefore applies to estimating societal value as well as outcome frequencies. In contrast, agents have subjective utilities over the outcomes that they wish to optimize, possibly different from the “societal value” or any other metric used by the RCT.
@@
We address participation incentives by imposing Bayesian Individual Rationality (BIR) constraints, a standard notion in economic theory. Specifically, the mechanism commits to some randomized policy that assigns an arm to each agent, depending on the history of interactions with the previous agents. The agent does not observe the history or the chosen arm (as discussed above). BIR states that each agent must prefer participation to her “outside option,” in expectation over her Bayesian beliefs.
@@
We consider three model variants. To isolate the core issues of incentivized participation, we first consider the fundamental case of homogeneous agents with the same beliefs and preferences. Then, the general case of heterogeneous agents endows each agent with a type. Then observable attributes such as medical history or user profile comprise one’s public type and determine one’s beliefs over outcomes. Moreover, one has unobservable preferences over the outcomes (private type) which need to be elicited in an incentive-compatible way. [^8] In practice, preference elicitation would happen during registration/opt-in process. For online platforms, private preferences and their elicitation make sense only if the RCT has explicit opt-in. The latter is uncommon, but may be needed if the RCT involves high-stakes interactions (such as buying a house) or is potentially controversial. The third model variant provides the mechanism with estimated type frequencies over the agent population. This final variant addresses situations in which some types are inherently (much) more “difficult” to handle than others. When such “difficult” types are rare, we leverage the type frequencies to mitigate their harmful influence.
Agents’ beliefs. While we allow adversarial outcomes, the agents believe that the outcome-generating process is much more well-behaved. Specifically, the agents believe that one’s outcome distribution is determined by the assigned arm and one’s public type (if any), independent of the round. We call such beliefs time-invariant. We posit that the beliefs are shared by all agents and are known to the mechanism, as an appropriately defined Bayesian prior.
This dichotomy between adversarial outcomes and simpler, time-invariant beliefs of the agents is essential to our model. On the one hand, our statistical guarantees are model-free once the data is collected, matching the standard for RCTs. On the other hand, time-invariant beliefs are standard in modeling sequential decision-makers (e.g., in incentivized exploration, see Related Work). While they need not be correct, they are reasonable as a model of reality and quite realistic as beliefs that patients or platform users might have, particularly so for clinical RCTs where they represent pre-existing medical knowledge. Beliefs whose support does not include the actual problem instance (a.k.a. misspecified beliefs) are not uncommon in economic theory, see Related Work.
Crucially, the BIR constraint only needs to hold with respect to agents’ time-invariant beliefs. Put differently, it suffices to provide participation incentives for the hypothetical world in which the agents’ beliefs are correct. We note that a stronger version of participation incentives — robust against an adversarially chosen outcome sequence — is impossible to achieve (cf. Footnote 16).
@@
Our results. We focus on mechanisms with simple two-stage designs: a relatively short warm-up stage of predetermined duration, followed by the main stage that needs to be counterfactually estimated. The data from the warm-up stage (warm-up data) creates information asymmetry in the main state which drives participation incentives, whereas the main stage determines the counterfactual estimates. RCTs with a warm-up stage are very common in practice: e.g., phase-I clinical trials to calibrate the dosage and “pilot experiments” in online platforms.
@@
Our guarantees focus on the main stage, given enough warm-up data. A relatively small amount of warm-up data, determined by the prior and the utility structure, suffices to bootstrap the main stage for an arbitrarily large number of agents. The warm-up data can be collected exogenously e.g., using paid volunteers, or endogenously by a mechanism in our model, e.g., by leveraging prior work on incentivized exploration (see “Related work” and Lemma 4.6).
@@
We obtain a near-optimal solution for each model variant described above, from homogeneous agents to heterogeneous agents to estimated type frequencies. First, we put forward an idealized benchmark which focuses on a single round of our problem. Then, we design a mechanism whose estimation error $\mathtt{ERR}$ (in our actual, multi-round model) is upper-bounded in terms of this single-round benchmark. Finally, we prove that no mechanism achieves better estimation error, up to constant factors. While the benchmark, the mechanism, and the lower bound are different for each model variant, many ideas carry over from one model variant to the next.
The single-round benchmark operates a hypothetical world in which the true “state of the world” is drawn according to the agents’ beliefs. The benchmark is a hypothetical policy that observes the true state (and the agent’s type, if applicable) and chooses a distribution over actions so as to optimize a certain objective: essentially, an upper bound on $\mathtt{ERR}$. Thus, the benchmark encapsulates the optimal dependence on the prior in this hypothetical scenario.
@@
Our mechanisms have a simple structure detailed below (which, however, suffices to connect them to the single-round benchmark). First, their “sampling distributions” — the distributions from which the arms are sampled — do not adapt to the data collected in the main stage. This is a desirable property: indeed, clinical RCTs in practice typically do not adapt to data, except possibly at a few pre-determined “review points”, [^9] This is because treating the patients and/or observing the outcomes may take a long time, and the doctors may find it difficult to coordinate across the patients. So, many/most patients may need to be treated in parallel. and neither do most RCTs in online platforms due to engineering considerations. Second, the type-dependent sampling distribution does not change throughout the main stage. In particular, for the homogeneous case we follow the epsilon-greedy paradigm from multi-armed bandits, choosing a random arm with fixed probability $\epsilon$ and an estimated best arm otherwise. [^10] The exploration probability $\epsilon$ is chosen in a non-standard way depending on the prior, and the estimated best arm is determined by the warm-up data. Third, counterfactual estimates are formed using the standard machinery of Inverse Propensity Scores (IPS). In contrast, our negative results apply to a wide class of mechanisms, allowing arbitrary dependence on the past data (in most rounds), arbitrary changes throughout the main stage, and arbitrary counterfactual estimates. The substance of our analysis is connecting both upper and lower bounds to the single-round benchmark.
@@
We emphasize that our contribution spans the whole “journey”: from the new model to the benchmark to the matching upper/lower bounds, and from homogeneous agents to heterogeneous agents to estimated type frequencies. Conceptually, we connect the dynamic process and statistical estimators in the mechanism with a static information-design problem in the benchmark. We focus on the asymptotically-optimal analysis of the main stage, given with some “constant” (prior-dependent) amount of warm-up data. While we characterize the sufficient amount of warm-up data, we suspect it could be improved.
@@
We focus on patients’ reluctance to participate that stems from their preferences among the available treatments (as opposed to other factors, e.g., logistical difficulties). Further, this paper is only concerned with patients’ initial decision to participate in an RCT. Their incentives to remain in the trial, e.g., if the treatment appears ineffective, are outside of our scope. [^11] This is indeed an issue in practice, as per (FDA,, 2001, Section 2.1.7.2). However, it is unclear if any approach based on information asymmetry could possibly help here: once the patient learns about the actual outcome, the “principal” does not have much/any informational advantage left.
Map of the paper. We first present and analyze our model for the case of homogeneous agents (Sections 3 and 4). This is to introduce the core model and ideas in a more accessible way, disentangling them from the considerable extra complexity (ideas, technicalities, and heavier notation) needed for the general case. The main case of heterogeneous agents is defined and treated in Section 5. The extension with estimated type frequencies is in Section 6. The technical machinery in Sections 4, 5 and 6 is introduced gradually and heavily reused from one section to the next.
2 Related work
Incentivized exploration (IE) is a closely related model concerning experimentation on online platforms. Here the “platform” is a recommendation system which strives to incentivize users to explore when they would prefer to exploit. The platform issues recommendations (which the users can override), and makes them incentive-compatible by leveraging information asymmetry and the power to commit to a particular policy. Introduced in (Kremer et al.,, 2014; Che and Hörner,, 2018), IE is well-studied, see (Slivkins,, 2019, Chapter 11) for a survey. Most technically related are Mansour et al., (2020); Immorlica et al., (2019).
The main difference concerns performance objectives. IE strives to maximize the total reward under a time-invariant outcome distribution, or just explore each arm once. Neither serves our statistical objective, which requires independent near-uniform randomization in each round: reward-maximization pulls a mechanism in a very different direction, and merely exploring each arm $N$ times does not ensure statistical validity under adversarially chosen outcomes. The economic constraints are different, too, because agents in our model can only choose whether to participate. [^12] This distinction is consequential only for $n>2$ arms. Then the economic constraint in IE is stronger than ours. Hence, our mechanisms, analyses and technical guarantees are very different from those in IE. However, our mechanisms reuse IE for the “warm-up stage” (see Lemma 4.6).
@@ 2 Related work
Economic design. Using information asymmetry and commitment power to incentivize agents to take a particular action is a common goal in information design (Bergemann and Morris,, 2019; Kamenica,, 2019). However, this literature tends to have a very different technical flavor. The key distinction is that the designer’s utility is (typically) directly derived from each agent’s actions. In contrast, our statistical objective depends on the sequence of sampling distributions and does not add up across agents.
Misspecified beliefs (i.e., beliefs whose support does not include the actual problem instance) are studied in a considerable recent literature in economic theory (starting from Berk,, 1966; Esponda and Pouzo,, 2016). However, this literature tends to focus on convergence and equilibria (e.g., Fudenberg et al.,, 2021; Esponda and Pouzo,, 2021; Esponda et al.,, 2021), whereas our paper takes the mechanism design perspective and is concerned with statistical estimation.
@@
Difficulties in patient recruitment are widely acknowledged as a significant barrier (e.g.,  Adams et al.,, 1997; Grill and Karlawish,, 2010). Its reasons are well-studied, e.g., see surveys (Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). Patients’ preferences among treatments are listed among the major factors, along with other factors such as insufficient trust, logistical difficulties, or aversion to randomization and loss of control. Prior work suggested increasing the probability of the “new” treatment as a possible remedy (Vozdolska et al.,, 2009; Jenkins and Fallowfield,, 2000; Pocock,, 1979; Karlawish et al.,, 2008), albeit without a formal model of participation incentives and without theoretical guarantees. These papers focus on the scenario when the new treatment appears ex-ante preferable compared to non-participation, and do not revise the treatment probabilities as the new data arrives. Other suggested remedies, focusing on eligibility criteria, patient outreach, trial logistics, etc., are not as relevant to this paper.
@@
Our mechanisms are superficially similar to patient preference trials (PPTs) (Torgerson and Sibbald,, 1998), in that both allow patient preferences to impact treatment allocation. The key difference is that PPTs allow the patients to directly choose the chosen treatment (potentially at the expense of internal validity), whereas in our setting a patient’s report only affects the sampling distribution (hence internal validity is not compromised). Another distinction is that we elicit patient preferences over the treatment outcomes, not over the treatments themselves.
@@
While our mechanism follows the protocol of multi-armed bandits (Slivkins,, 2019; Lattimore and Szepesvári,, 2020), it strives to choose arms uniformly at random, whereas a bandit algorithm strives to converge on the best arm. [^13] Under the standard objective of cumulative reward/loss optimization, as expressed by regret. More related is “pure exploration” (starting from Mannor and Tsitsiklis,, 2004; Even-Dar et al.,, 2006), where a bandit algorithm predicts the best arm under a stationary outcome distribution, after exploring “for free” (e.g., uniformly) for $T$ rounds. In contrast, our model allows adversarially chosen outcomes and requires estimates for all arms and all outcomes. Note that bandit algorithms, whether regret-minimizing or “pure exploration”, are generally not geared to incentivize participation. Interestingly, both “stationary” and adversarially chosen outcome distributions – like, resp., agent beliefs and statistical objective in our model – are well-studied in bandits, as, resp., “stochastic” and “adversarial” bandits.